OpenStudy (anonymous):

a 5.49g sample of iron(iii) chloride hydrate was found to contain 2.20g of water. what is the complete formula of the hydrate?

OpenStudy (anonymous):

%(H2O)=2.2/5.49 x 100 =40% so m(FeCl3) = 60% m(total) M(FeCl3)= 162.2g M(FeCl3. xH2O)= 162.2 x100/60 =270.3g so m(H2O) = 270.3-162.6 = 108.13g n(H2O) =m/M =108.13/ (2x1.008+16) 6.002 mol (in one mol of FeCl3) thus hydrate formula= FeCl3. 6H2O